OpenStudy (anonymous):

If you roll a number cube 12 times, about how many times would you expect to roll a 5 or 6? A. 2 B. 5 C. 4 D. 10

OpenStudy (anonymous):

\[P(5 or 6)=P(5)+P(6)-P(5 and 6)=P(5)+P(6)\]\[=\frac{1}{6}+\frac{1}{6}=\frac{1}{3}\]

OpenStudy (anonymous):

So 1/3 of 12 times you'd expect to get a 5 or a 6; i.e. 4 times.